# Semblance study of force-field and convection-diffusion solutions with observations of solar magnetic phenomena

### v2

@@ 7 Count rates from ground-based detectors
The neutrons detected at ground level are the product of atmospheric showers caused by the collision of a primary particle, that collides with a molecule in the atmosphere. The neutron count is given by the following equation:
@@
Where $j_{i}(P,t)$ is the spectrum of the primary species above the atmosphere, and $S_{i}(P,x)$ is the atmospheric yield function due to this species. $P_{c}$ is the cut-off rigidity, which is the necessary minimum energy for a particle to enter the Earth’s magnetosphere. The quantity $dN=dP$ is the differential counting rate of the instrument inside the atmosphere. The yield function, that was used in this analysis is the following (Caballero-Lopez & Moraal, 2012):
@@
where $F(P)$ is the ratio between the yield functions of $He$ and $H$ reported by (Clem & Dorman, 2000):
@@
The equation 42 allows to compute a modulation factor and will be used to obtain the semblance. The $J_{H}$ and $J_{He}$ terms in the equation 42 are the transport equation solution (convection-diffusion or force field) that are in terms of a LIS previously mentioned. A code on Python has been programmed and can be found here: https://github.com/JorUrie/TRC.git
@@ 8 First Analysis
The first analysis consists in three stations located in different latitudes (Table 2) and compute their modulation factor, then semblance between modulation factor with sunspots or mean solar magnetic field will be computed. Data were downloaded from the Neutron Monitor DataBase (NMDB) [^3] NMDB, sunspots data from Sunspot Number [^4] Sunspot Number | SILSO (sidc.be) (Figure 3) and mean solar magnetic field from The Wilcox Solar Observatory [^5] WSO - The Wilcox Solar Observatory (stanford.edu) (Figure 4).
@@
SNAE 0.73
OULU 0.81
AATB 5.90
@@
The Figure 3 is about the number of sunspots recorded, that are related to solar maximum and minimum. Data is from May, 1957 to July, 2022. In the same figure, solar cycles can be observed, in other words, solar maximums belong to the years 1958, 1970, 1980, 1990, 2014 and the next solar maximum could be in 2024. Solar minimum belongs to the years 1965, 1976, 1986, 1997 and 2010. With the previous information, 11-year cycle can be perceived. In 1958 was the larggest recorded maximum, while the smallest maximum was in 1970, as soon as minimums have similar values and there are not much differences.
Mean magnetic field information belongs to the dates from May 1975 to December 2015. The information has cusps and looks to not have changes, but in September 1981 was the biggest registered maximum, while in March 1991 was the smaller registered value. If a line is traced along the graphic following a ”regression”, maximums and minimums could be seen, they would be mostly inversely proportionate to sunspots.
The selected stations were SNAE (south), OULU (north) and AATB (equator). They were deemed appropriate based on their respective latitudes in Earth (Figure 24). SNAE, OULU and AATB data are showed in the figures 5, 6 and 7. Data were downloaded and normalized by contained higher value.
@@
The previous mentioned figures 5, 6 and 7 are normalized data, because the data is obtained from the neutral monitors on the Earth, however the particles have already modified by modulation phenomena and they have also lost energy (secondary cosmic rays), for that reason a modulation factor is necessary to compute.
Non-normalised data have been introduced into the atmospheric yield function, which requires a LIS and a transport equation solution (convection-diffusion or force field) to compute the modulation factor.
The figures from 8 to 15 are the modulation factors for the same station with every previously mentioned LIS and a transport equation solution. Their unities are $Counts/s$ and $Date$, because the modulation factor is the amount of particles before to enter the atmosphere and be influenced by particles or the magnetosphere of the Earth (primary species) in a given date. Modulation factor would not be necessary if the satellites would take desired measurements, but the technology and the conditions do not allow it.
@@
The previous compute was applied to all selected stations. The spectral modulation factors have a similar pattern for each LIS and are also similar to the downloaded data, for example, the figure 6 and figure 8, can be observed the high values in 1987, near to 2004 and 2011, those occurred for the figures 8 to 15, in fact for the figures 8, 10, 12 and 14 between values $1.2$ and $1.4$ $Counts/s$, two points are always present and are also the highest values. For the figures 9, 11, 13 and 15, they have the same pattern and the two highest points are also present, but between $1.5$ and $1.75$ $Counts/s$, so the difference of the spectral modulation factors using convection-diffusion solution and force field solution is the amplitude.
From the figures 8 to 15 also looks like to the figure 3, because there are more extragalatic cosmic rays introduced into the Earth as a result of the low solar modulation.
For the three stations, that have been selected according to their latitude, the semblance between modulation factor and sunspots using several LIS are showed, however only for OULU station, the rest models can be found in the appendix.
@@
The computed models offer similar results among them (figures $a)$ and $b)$ for each semblance figure), they even look like if they were the same model, but thanks to the difference between semblances (figures $c)$) is possible to observe, that there are little differences mainly in the figures 21 and 20. If both figures previously mentioned are comparative with the rest of respective semblances in this section, the main difference is in the interval 50 to 100 in horizontal axis and 40 to 80 in the wavelength, where blue spot change his size and form.
There is correlation between sunspots and modulation phenomena, which can be observed in the figures 16, 18, 20 and 22 highlighted in red. There is anti-correlation between the mean solar magnetic field and the modulation phenomena observed in blue colour. Both behaviours are maintained independently of the Earth’s latitude.
@@ 9 Second Analysis
The same calculus was performed, but the spectral factor modulations were average while taking into account the LIS. The data from the following stations were used: KERG, MOSC, OULU, THUL, and HRMS (Table 3).
@@
THUL 0.3
OULU 0.81
KERG 1.14
MOSC 2.43
HRMS 4.58
@@
The Figures 25 to 32 are the spectral averaged using convection-diffusion and force field solutions with Lagner, Potgieter & Webber LIS in 2003, Burguer & Potgieter LIS in 2000, Garcia-Munoz, Mason & Simpson LIS in 1975 and Ghelfi, Barao, Derome & Maurin LIS in 2017.
@@
The Figures 33 to 48 are the semblance between the aforementioned modulation factors with sunspots and mean magnetic field data. The figures 34, 36, 38, 40 and 42 are the percentage of contained zeros into the difference between semblance (Figures $c)$), that were calculated to observed with more detail which LIS has the best adjusted in order to find the most reliable model.
@@
In this case, an average modulation factor of several neutron detectors around the word was computed to analyse the general behaviour between sunspots and mean solar magnetic field with modulation phenomena. The results were utterly similar to the first analysis, in fact the analysis of the contained zeros confirms the hypothesis, that there is correlation between sunspots and modulation phenomena and anti-correlation between the mean magnetic field and modulation phenomena. The analysis of the zeros can be summarised in the Table 4.
@@ 10 Conclusions
Both analyses allow quantifying the correlation between modulation factor with sunspots and anti-correlation between modulation factor with mean magnetic field. The models are mostly similar with both transport equation solutions and the local interstellar spectrum, because just shown light changes.
For the first analysis, the modulation factors yield similar outcomes and are solely dependent on the rigidity factor. The correlation between sunspots and modulation factor is more pronounced for stations closer to the poles, such as OULU and SNAE, than the equatorial AATB station. However, this pattern is not replicated for the semblance between the modulation factor and the mean magnetic field. The station near the equator shown more anti-correlation than near the poles. The pattern is similar for every semblances, which were used different LIS and transport equation solution.
Thanks to the second analysis, the first analysis could be affirmed and generalized, the force field solution is also more compatible with neutron detectors data, Additionally any modulation factor solution will be similar to the Neutron detector data and just will have minimal amplitude variations.
In the second analysis, it is possible to observe similarity among the graphs of contained zeros into the difference of semblances, particularly when the modulation factors have been computed by different LIS. However if semblance is computed using the Garcia-Munoz, Mason and Simpson LIS in 1975 shown more differences, that is visible in the analysis of contained zeros for semblance between averaged modulation factor and mean magnetic field.
According to the table 4 semblance between sunspots and averaged modulation factors are really similar and the percentage of contained zeros are akin if Webber & Lockwood and Burguer and Potgieter LIS in 2000 are used, same happened for Garcia-Munoz and Ghelfi, Barao, Derome and Maurin LIS in 2017.
Therefore, the LIS that have the best adjust to the semblance model are Lagner, Potgieter and Webber LIS in 2003, Burguer and Ghelfi, Barao, Derome and Maurin LIS in 2017 using force field solution because they offer more consistent values as shown in the table 4.
With both proposed models can be asserted with greater certainty that exist a correlation between modulation phenomena and sunspots, because the solar activity is less allowed to detect more rays from the Sun. The anti-correlation between modulation phenomena and mean magnetic field is due to the high solar activity, therefore there are less solar rays that affect the arisen of cosmic rays from others source.
Thanks to this proposed quantitative model, a more precise predictive model of cosmic rays with solar activity can be compute.

### v3

@@ 7 Count rates from ground-based detectors
The neutrons detected at ground level are the product of atmospheric showers caused by the collision of primary particles with molecules in the atmosphere. The neutron count is given by the following equation:
@@
Where $j_{i}(P,t)$ is the spectrum of the primary species above the atmosphere, and $S_{i}(P,x)$ is the atmospheric yield function due to this species. $P_{c}$ is the cut-off rigidity, which is the necessary minimum energy for a particle to enter the Earth’s magnetosphere. The quantity $dN=dP$ is the differential counting rate of the instrument inside the atmosphere. The yield function used in this analysis is the following (Caballero-Lopez & Moraal, 2012):
@@
where $F(P)$ is the ratio between the yield functions of $He$ and $H$ as reported by (Clem & Dorman, 2000):
@@
The equation 42 allows computing a modulation factor and will be used to obtain the semblance. The $J_{H}$ and $J_{He}$ terms in the equation 42 are the transport equation solutions (convection-diffusion or force field), which are in terms of a LIS previously mentioned. A Python code has been programmed and can be found at https://github.com/JorUrie/TRC.git
@@ 8 First Analysis
The first analysis consists of three stations located in different magnetospheric latitudes (Table 2) and and computing their modulation factor, then semblance between the modulation factor and sunspots or mean solar magnetic field will be computed. The data were downloaded from the Neutron Monitor DataBase (NMDB) [^3] NMDB, sunspot data from Sunspot Number [^4] Sunspot Number | SILSO (sidc.be) (Figure 3) and mean solar magnetic field data from The Wilcox Solar Observatory [^5] WSO -The Wilcox Solar Observatory (stanford.edu) (Figure 4).
@@
Antarctic (SNAE) 0.73
Finland (OULU) 0.81
Kazakhstan (AATB) 5.90
@@
The Figure 3 shows the number of sunspots recorded, which are related to solar maximums and minimums. The data covers from May 1957 to July 2022. In the same figure, solar cycles can be observed, in other words, solar maximums belong approximately to the years 1958, 1970, 1980, 1990, 2014 and the next solar maximum could be in 2024. Solar minimums belong approximately to the years 1965, 1976, 1986, 1997 and 2010. With the previous information, 11-year cycle can be perceived. In 1958 was the largest recorded maximum, while the smallest maximum was in 1970. Conversely, minimums have similar values and there are not many differences.
The mean magnetic field information belongs to the dates from May 1975 to December 2015. The information has cusps and appears not to have significant changes, but in September 1981 was the biggest registered maximum, while in March 1991 was the smallest registered value. If a line is traced along the graphic following a “regression”, maximums and minimums can be identified, they would be mostly inversely proportionate to sunspots.
The selected stations were SNAE (south), OULU (north) and AATB (magnetic equator). They were deemed appropriate based on their respective latitudes in the Earth (Figure 24). SNAE, OULU and AATB data are shown in the figures 5, 6 and 7. The data were downloaded and normalized to the highest contained value.
@@
The previously mentioned Figures 5, 6 and 7 are normalized data, because the data is obtained from the neutron monitors on the Earth. However, the particles have already been modified by modulation phenomena and they have also lost energy (secondary cosmic rays), for that reason a modulation factor is necessary to compute.
Non-normalized data have been introduced into the atmospheric yield function, which requires both a LIS and a transport equation solution (convection-diffusion or force field) to compute the modulation factor.
The figures from 8 to 15 are the modulation factors for the same station using every previously mentioned LIS and a transport equation solution. Their unities are $Counts/s$ and $Date$, because the modulation factor is the amount of particles before they enter the Earth’s atmosphere and interact with the particles in a given date. Modulation factor would not be necessary if the satellites would take desired measurements, but the technology and the conditions do not allow it.
@@
The modulation factors were computed for all selected stations. Those have a similar pattern for each LIS and are also similar to the downloaded data, for example, in Figures 6 and 8, high values can be observed, in 1987 near to 2004 and 2011, These patterns also occur in Figures 8 to 15. In fact, in Figures 8, 10, 12 and 14 between values $1.2$ and $1.4$ $Counts/s$, two points are always present and are also the highest values. For the figures 9, 11, 13 and 15, the same pattern is observed and the two highest points are also present between $1.5$ and $1.75$ $Counts/s$. The main difference among modulation factors using convection-diffusion and force field solutions is the amplitude.
From Figures 8 to 15 also looks like the figure 3, because more extragalactic cosmic rays are introduced into the Earth as a result of low solar modulation.
For the three stations selected according to their latitude, the semblance for the modulation factor and sunspots using several LIS are shown. However, only the results for the OULU station are presented here; the remaining models can be found in the appendix.
@@
The computed models offer similar results among them (figures $a)$ and $b)$ for each semblance figure). They even appear as if they were the same model, but thanks to the difference between semblances (figures $c)$) it is possible to observe small differences, mainly in Figures 21 and 20. If both figures previously mentioned are compared with the rest of the respective semblances in this section, the main difference is in the interval 50 to 100 in number of data and 40 to 80 in the wavelength, where a blue spot changes its size and shape.
There is correlation between sunspots and modulation phenomena, which can be observed in the figures 16, 18, 20 and 22 highlighted in red. There is anti-correlation between the mean solar magnetic field and the modulation phenomena observed in blue. Both behaviors are maintained independently of the magnetospheric latitude.
@@ 9 Second Analysis
The same calculations were performed, but the modulation factors were averaged. The data used for the following analysis were: KERG, MOSC, OULU, THUL, and HRMS (Table 3). This analysis was conducted to generalize the previous results and to analyze which LIS have the best fit.
@@
Greenland (THUL) 0.3
Finland (OULU) 0.81
Kerguelen (KERG) 1.14
Moscow (MOSC) 2.43
Hermanus (HRMS) 4.58
@@
The Figures 25 to 32 are the modulation factors averaged using convection-diffusion and force field solutions with Lagner, Potgieter & Webber LIS in 2003, Burguer & Potgieter LIS in 2000, Garcia-Munoz, Mason & Simpson LIS in 1975 and Ghelfi, Barao, Derome & Maurin LIS in 2017.
@@
The Figures 33 to 48 show the semblance between the aforementioned modulation factors with sunspots and mean magnetic field data. The figures 34, 36, 38, 40 and 42 show the percentage of zeros contained in the difference between semblance (Figures $c)$). Thanks to this analysis, the LIS models were selected because they offer the most reliable results.
@@
In this case, an average modulation factor of several neutron detectors around the world was computed to analyze the general analyze between sunspots and mean solar magnetic field and modulation phenomena. The results were utterly similar to the first analysis, in fact the analysis of the contained zeros confirms the hypothesis, that there is a correlation between sunspots and modulation phenomena and an anti-correlation between the mean magnetic field and modulation phenomena. The analysis of the zeros is summarized in Table 4.
@@ 10 Conclusions
Both analyses allow quantifying the correlation between the modulation factor and sunspots and anti-correlation between the modulation factor and the mean magnetic field. The models are mostly similar with both transport equation solutions and the local interstellar spectrum as they only show slight changes.
For the first analysis, the the modulation factor yields similar outcomes and are solely dependent on the rigidity factor. The correlation between sunspots and modulation factor is more pronounced for stations closer to the magnetic poles, such as OULU and SNAE, than for the equatorial AATB station. However, this pattern is not replicated for the semblance between the modulation factor and the mean magnetic field. The station near the magnetic equator has shows more anti-correlation than near the poles. The pattern is similar for every semblance, even when different LIS and transport equation solutions were used.
Thanks to the second analysis, the first analysis can be affirmed and generalized, the force field solution is also more compatible with neutron detectors data. Additionally, any modulation factor solution will be similar to the neutron detector data and will only have minimal amplitude variations.
In the second analysis, it is possible to observe similarity among the graphs of contained zeros in the difference of semblances, particularly when the modulation factors have been computed by different LIS. However, if semblance is computed using the Garcia-Munoz, Mason and Simpson LIS in 1975 more differences are observed. These differences are visible in the analysis of contained zeros for semblance between averaged modulation factor and mean magnetic field.
According to the table 4 semblance using sunspots and averaged modulation factors data is very similar, and the percentage of contained zeros is comparable when using the Webber & Lockwood and Burguer and Potgieter LIS in 2000 or Garcia-Munoz and Ghelfi, Barao, Derome and Maurin LIS in 2017 are used.
Therefore, the LIS, that have the best adjust to the semblance model are Lagner, Potgieter and Webber LIS in 2003, Burguer and Ghelfi, Barao, Derome and Maurin LIS in 2017 using force field solution as they offer more consistent values, as shown in Table 4.
With both proposed models, it can be asserted that there is a correlation between modulation phenomena and sunspots, because the solar activity is less allowed to detect more rays from the Sun. The anti-correlation between modulation phenomena and the mean magnetic field is due to high solar activity. Therefore, fewer solar rays affect the arrival of cosmic rays from other sources.
Thanks to this proposed quantitative model, a more precise predictive model of cosmic rays with solar activity can be computed.